CS 582 — Database Management System II— Spring 2007

Homework 3 Deadline: 2/15/2007 – 11:55pm

The purpose of this homework is to refresh our PHP skill. Write a PHP program that accesses your database (created for homework 2) and displays the following:

• List of all students with their transcripts (student should be listed with their id and name; course name should be listed instead of course code, for each student the cumulative GPA should be computed);

• List all professors with their departments.

Submit your PHP as the answer for this assignement. This will be kept for the record. Send an email to me to indicate your PHP website so your work can be graded.)

Note: I have a few programs that you can use for accessing Oracle database from PHP that is located at

http://www.cs.nmsu.edu/~tson/classes/materials/PHP/